# StableToken: A Noise-Robust Semantic Speech Tokenizer for Resilient SpeechLLMs

### v1

@@ Abstract
Prevalent semantic speech tokenizers, designed to capture linguistic content, are surprisingly fragile. We find they are not robust to meaning-irrelevant acoustic perturbations; even at high Signal-to-Noise Ratios (SNRs) where speech is perfectly intelligible, their output token sequences can change drastically, increasing the learning burden for downstream LLMs. This instability stems from two flaws: a brittle single-path quantization architecture and a distant training signal indifferent to intermediate token stability. To address this, we introduce StableToken, a tokenizer that achieves stability through a consensus-driven mechanism. Its multi-branch architecture processes audio in parallel, and these representations are merged via a powerful bit-wise voting mechanism to form a single, stable token sequence. StableToken sets a new state-of-the-art in token stability, drastically reducing Unit Edit Distance (UED) under diverse noise conditions. This foundational stability translates directly to downstream benefits, significantly improving the robustness of SpeechLLMs on a variety of tasks.
@@ 1 Introduction
The application of Large Language Models (LLMs) to the speech domain has given rise to a new class of powerful models: Speech Large Language Models (SpeechLLMs) (Hurst et al., 2024; Défossez et al., 2024; Zeng et al., 2024). These models rely on a discrete speech tokenizer to convert continuous audio into tokens sequences that the LLM can process. Among available methods, semantic tokenizers have been widely adopted, as their low-bitrate, semantically-aligned outputs are highly compatible with LLM architectures (Défossez et al., 2024; Zeng et al., 2024; Ding et al., 2025; Wu et al., 2025).
@@
Despite their widespread adoption and apparent success, we find that these semantic tokenizers harbor a critical vulnerability: a profound lack of robustness. Contrary to their core design principle of encoding semantics, even imperceptible acoustic noise can induce drastic shifts in their discrete outputs (Figure 1). The general issue of tokenizer instability is also supported by recent findings on earlier non-VQ-based SSL tokenizers (Messica & Adi, 2024). This instability creates a damaging downstream effect: small acoustic changes trigger large token jumps, which break the crucial speech–text alignment and pose a significant modeling challenge for the LLM, forcing it to learn from an inconsistent or even chaotic input stream. This inherent fragility is severely amplified by environmental noise, serving as a key cause for the performance degradation of SpeechLLMs in real-world conditions (Ma et al., 2025b; Zhang et al., 2025; Yang et al., 2024c; Jiang et al., 2025). We argue that enhancing tokenizer robustness is therefore a direct and promising path toward building more resilient models.
@@
Addressing this dual challenge is non-trivial. For the brittle architecture, an offline ensemble of models seems intuitive. However, this approach is untenable: (1) it prohibitively increases inference cost; (2) aggregating independently trained models is non-trivial, as their quantization boundaries are arbitrarily aligned; and (3) a token-level majority vote is too coarse. To tackle the distant supervisory signal, one might introduce a token-level consistency objective for clean and noisy input audios. Yet, this leads to notoriously unstable gradients when applied to discrete codes, making the model difficult to train. The failure of these straightforward approaches underscores the need for a more integrated paradigm.
@@ 2.1 Overall Structure
Our approach, StableToken, is designed to overcome the fragility of prevailing VQ-based semantic tokenizers. These tokenizers often produce unstable token sequences in the presence of subtle noise, a vulnerability stemming from two core weaknesses: (1) a single-path architecture that lacks fault tolerance, and (2) a distant supervisory signal that fail to enforce representational invariance.
@@ 2.2 The Voting-LFQ Module
By using an odd number of parallel branches ($n$), we enforce a strict majority rule via a bit-wise vote, creating exceptional robustness against noise. This approach not only corrects errors when a minority of branches fail, but can also recover the true token even if a majority of branches are corrupted at the token-level. Recovery is possible as long as the underlying bit-level errors remain sparse. This resilience marks a significant advantage over fragile single-path quantizers and, in parallel, allows for more expressive representations during training.
@@ 3 Experimental Setup
We benchmark StableToken against a comprehensive suite of SOTA models across three categories: SSL-based, distilled, and supervised tokenizers. For all baselines, we use their officially released models. The detailed list of baseline models can be found in Appendix C.
@@
We assess the tokenizer’s intrinsic properties. Robustness is measured by Unit Edit Distance (UED%, $\downarrow$(Messica & Adi, 2024) on the FLEURS (Conneau et al., 2023) benchmark under various synthetic perturbations and real-world noise conditions, which include challenging out-of-domain (OOD) noise. A detailed description of the noise profiles is in Appendix D. Fidelity is measured by Word Error Rate (WER%, $\downarrow$) and Mean Opinion Score (MOS, $\uparrow$) on the LibriSpeech (Panayotov et al., 2015) and SEED (Anastassiou et al., 2024) benchmarks.
@@
NAST (Messica & Adi, 2024) 1 50Hz 200 18.67 15.78 15.26 14.95 18.69 19.07 17.07
R-Spin (Chang & Glass, 2024) 1 50Hz 2048 21.56 17.08 15.47 14.95 15.08 14.75 16.48
@@
For downstream tasks, we follow a controlled, isogenic setup to ensure fair comparison. Each tokenizer is integrated into a SpeechLLM framework using a pre-trained Qwen2.5-3B (Yang et al., 2024a) backbone, which is then fine-tuned using a prompt-based paradigm (Zeng et al., 2025). We evaluate on three tasks: (1) ASR: Assessed on noise-augmented LibriSpeech (Panayotov et al., 2015) and the CHiME-4 (Vincent et al., 2017) benchmark using WER (%); (2) SER: Assessed on a noise-augmented version of the ESD (Zhou et al., 2022) test set using classification accuracy (%); (3) TTS: Assessed on the SEED-TTS (Anastassiou et al., 2024) benchmark using both WER (%) and MOS. The aggregated training datasets, fine-tuning hyperparameters, and prompts for each task are detailed in Appendix E.
@@ 4.3.3 Case Study
Table 6 provides a case study illustrating the error correction capability of the voting-LFQ module. For instance, at position 80, noise causes three voters to generate erroneous tokens. Specifically, Voters 2 and 3 flip bit #5, while Voter 5 flips bit #7. Despite most voters predicting incorrect tokens, the voting mechanism operating at the bit level allows for correct recovery. For bit #5, the correct value ’0’ wins by a 3-to-2 majority, and for bit #7, the correct value ’1’ wins by a 4-to-1 majority, successfully reconstructing the original token (3485). Similar corrections occur at positions 68, 105 and 114. This case study highlights a key advantage of StableToken: its resilience does not depend on every branch being perfect, but on the collective ability to override sparse bit-flip errors.
@@ 5 Related Work
SSL tokenizers rely on self-supervised learning (SSL) to extract discrete units from audio (Chen et al., 2022; Chung et al., 2021; Conneau et al., 2021; Chiu et al., 2022), but are proven suboptimal for generative speechLLMs (Mousavi et al., 2024; Guo et al., 2025). Hybrid approaches combine acoustic tokenizers with semantic distillation to balance fidelity and content (Zhang et al., 2023; Ye et al., 2025; Siahkoohi et al., 2022; Yang et al., 2024b), yet their high data rate and structural incompatibility with LLMs introduce practical challenges. Recent fully supervised methods have proven suitable for SpeechLLMs, which produce linguistic tokens while retaining sufficient phonetic information (Zeng et al., 2025; Du et al., 2024a), supporting expressive and efficient end-to-end SpeechLLMs (Zeng et al., 2024; Ding et al., 2025; Wu et al., 2025).
@@
While extensive research has focused on constructing robust Automatic Speech Recognition (ASR) models (Wang et al., 2022; Tjandra et al., 2023; Eickhoff et al., 2023; Gong et al., 2023; Ahn et al., 2025), the stability of discrete speech tokens under noisy conditions has received much less attention. Recently, R-SPIN (Chang & Glass, 2024) and NAST (Messica & Adi, 2024) have begun addressing this gap, but are limited to traditional SSL-based speech tokenizers.
A more detailed discussion of related work is presented in Appendix G due to page limit.
@@ 6 Conclusion
To facilitate reproducibility of our work, we have provided detailed descriptions of the datasets, hyperparameters, and other experimental details used in our study in Section 3 and Appendix B, D, E. Our code and model checkpoint will be released publicly upon acceptance to further support reproducibility and foster future research.
@@ B.1 Training Datasets for StableToken
VCTK (Veaux et al., 2017) 44 ASR English
@@
VoiceAssistant-400K (Xie & Wu, 2024) 679 ASR English
AISHELLL-1 (Bu et al., 2017) 150 ASR Chinese
@@ B.5 Audio Reconstruction Details
Following the framework of CosyVoice (Du et al., 2024a) and GLM-4-Voice (Zeng et al., 2025), we train a flow matching model to reconstruct audio from speech tokens. The model takes as input the speech token representations and produces Mel spectrograms. Finally, a HiFi-GAN vocoder (Kong et al., 2020) converts the generated Mel spectrograms into the speech waveforms. The datasets for training the flow matching model is listed in Table 10 (excluding our in-house datasets, which comprise both English and Chinese speech data), and the training hyperparameters in Table 11.
B.6 Complexity Analysis of Different Voter Counts
@@
Table 12: Model parameters and inference FLOPs for different voter counts NN.
@@
The inference latency introduced by increasing the number of voters is also negligible. This is because all voter modules can be computed in parallel, and each voter processes the same input features. Furthermore, the computation within each voter is lightweight compared to the model backbone. Therefore, the overall inference time is dominated by the backbone, and increasing the number of voters does not noticeably affect the total latency.
Appendix C Baseline Models

### v2

@@ Abstract
Prevalent semantic speech tokenizers, designed to capture linguistic content, are surprisingly fragile. We find they are not robust to meaning-irrelevant acoustic perturbations; even at high Signal-to-Noise Ratios (SNRs) where speech is perfectly intelligible, their output token sequences can change drastically, increasing the learning burden for downstream LLMs. This instability stems from two flaws: a brittle single-path quantization architecture and a distant training signal indifferent to intermediate token stability. To address this, we introduce StableToken, a tokenizer that achieves stability through a consensus-driven mechanism. Its multi-branch architecture processes audio in parallel, and these representations are merged via a powerful bit-wise voting mechanism to form a single, stable token sequence. StableToken sets a new state-of-the-art in token stability, drastically reducing Unit Edit Distance (UED) under diverse noise conditions. This foundational stability translates directly to downstream benefits, significantly improving the robustness of SpeechLLMs on a variety of tasks. Our code and model are publicly available at https://github.com/Tencent/StableToken.
@@ 1 Introduction
The application of Large Language Models (LLMs) to the speech domain has given rise to a new class of powerful models: Speech Large Language Models (SpeechLLMs) (Hurst et al., 2024; Défossez et al., 2024; Zeng et al., 2024). These models rely on a discrete speech tokenizer to convert continuous audio into token sequences that the LLM can process. Among available methods, semantic tokenizers have been widely adopted, as their low-bitrate, semantically-aligned outputs are highly compatible with LLM architectures (Défossez et al., 2024; Zeng et al., 2024; Ding et al., 2025; Wu et al., 2025).
@@
Despite their widespread adoption and apparent success, we find that these semantic tokenizers harbor a critical vulnerability: a profound lack of robustness. Contrary to their core design principle of encoding semantics, even imperceptible acoustic noise can induce drastic shifts in their discrete outputs (Figure 1). The general issue of tokenizer instability is also supported by recent findings on earlier non-VQ-based SSL tokenizers (Messica and Adi, 2024). This instability creates a damaging downstream effect: small acoustic changes trigger large token jumps, which break the crucial speech–text alignment and pose a significant modeling challenge for the LLM, forcing it to learn from an inconsistent or even chaotic input stream. This inherent fragility is severely amplified by environmental noise, serving as a key cause for the performance degradation of SpeechLLMs in real-world conditions (Ma et al., 2025b; Zhang et al., 2025; Yang et al., 2024c; Jiang et al., 2025). We argue that enhancing tokenizer robustness is therefore a direct and promising path toward building more resilient models.
@@
Addressing this dual challenge is non-trivial. For the brittle architecture, an offline ensemble of models seems intuitive. However, this approach is untenable: (1) it prohibitively increases inference cost; (2) aggregating independently trained models is non-trivial, as their quantization boundaries are arbitrarily aligned; and (3) a token-level majority vote is too coarse. To tackle the distant supervisory signal, one might introduce a token-level consistency objective for clean and noisy input audios. Yet, this leads to unstable gradients when applied to discrete codes, making the model difficult to train. The failure of these straightforward approaches underscores the need for a more integrated paradigm.
@@ 2.1 Overall Structure
Our approach, StableToken, is designed to overcome the fragility of prevailing VQ-based semantic tokenizers. These tokenizers often produce unstable token sequences in the presence of subtle noise, a vulnerability stemming from two core weaknesses: (1) a single-path architecture that lacks fault tolerance, and (2) a distant supervisory signal that fails to enforce representational invariance.
@@ 2.2 The Voting-LFQ Module
By using an odd number of parallel branches ($n$), we enforce a strict majority rule via a bit-wise vote, creating exceptional robustness against noise. This approach not only corrects errors when a minority of branches fail, but can also recover the true token even if a majority of branches are corrupted at the token level. Recovery is possible as long as the underlying bit-level errors remain sparse. This resilience marks a significant advantage over fragile single-path quantizers and, in parallel, allows for more expressive representations during training.
@@ 3 Experimental Setup
We benchmark StableToken against a comprehensive suite of SOTA models across three categories: SSL-based, distilled, and supervised tokenizers. For all baselines, we use their officially released models. The detailed list of baseline models can be found in Appendix D.
@@
We assess the tokenizer’s intrinsic properties. Robustness is measured by Unit Edit Distance (UED%, $\downarrow$(Messica and Adi, 2024) on the FLEURS (Conneau et al., 2023) benchmark under various synthetic perturbations and real-world noise conditions, which include challenging out-of-domain (OOD) noise. A detailed description of the noise profiles is in Appendix E. Fidelity is measured by Word Error Rate (WER%, $\downarrow$) and Mean Opinion Score (MOS, $\uparrow$) on the LibriSpeech (Panayotov et al., 2015) and SEED (Anastassiou et al., 2024) benchmarks.
@@
NAST (Messica and Adi, 2024) 1 50Hz 200 18.67 15.78 15.26 14.95 18.69 19.07 17.07
R-Spin (Chang and Glass, 2024) 1 50Hz 2048 21.56 17.08 15.47 14.95 15.08 14.75 16.48
@@
For downstream tasks, we follow a controlled, isogenic setup to ensure fair comparison. Each tokenizer is integrated into a SpeechLLM framework using a pre-trained Qwen2.5-3B (Yang et al., 2024a) backbone, which is then fine-tuned using a prompt-based paradigm (Zeng et al., 2025). We evaluate on three tasks: (1) ASR: Assessed on noise-augmented LibriSpeech (Panayotov et al., 2015) and the CHiME-4 (Vincent et al., 2017) benchmark using WER (%); (2) SER: Assessed on a noise-augmented version of the ESD (Zhou et al., 2022) test set using classification accuracy (%); (3) TTS: Assessed on the SEED-TTS (Anastassiou et al., 2024) benchmark using both WER (%) and MOS. The aggregated training datasets, fine-tuning hyperparameters, and prompts for each task are detailed in Appendix F.
@@ 4.3.3 Case Study
Table 6 provides a case study illustrating the error correction capability of the voting-LFQ module. For instance, at position 80, noise causes three voters to generate erroneous tokens. Specifically, Voters 2 and 3 flip bit #5, while Voter 5 flips bit #7. Despite most voters predicting incorrect tokens, the voting mechanism operating at the bit level allows for correct recovery. For bit #5, the correct value ‘0’ wins by a 3-to-2 majority, and for bit #7, the correct value ‘1’ wins by a 4-to-1 majority, successfully reconstructing the original token (3485). Similar corrections occur at positions 68, 105 and 114. This case study highlights a key advantage of StableToken: its resilience does not depend on every branch being perfect, but on the collective ability to override sparse bit-flip errors.
@@ 5 Related Work
The evolution of LLMs has driven the transition of spoken dialogue models from traditional pipelines to end-to-end SpeechLLMs (Zhang and Wang, 2019; Zhang et al., 2020; Jacqmin et al., 2022; Lee et al., 2021; Fang et al., 2024; Défossez et al., 2024; Wang et al., 2024), with semantic tokenizers becoming increasingly crucial. The design of semantic tokenizers has evolved through several distinct paradigms. Early approaches utilized self-supervised learning (SSL) to derive discrete units from unlabeled data (Hsu et al., 2021; Baevski et al., 2020; Chen et al., 2022; Chung et al., 2021; Conneau et al., 2021; Chiu et al., 2022; Baevski et al., 2019; Liu et al., 2023; Gat et al., 2023; Huang et al., 2022; Lodagala et al., 2023; Chang et al., 2023). The vast majority of tokens produced by these methods are designed for discriminative tasks. It is reported that discretized SSL tokens primarily encode phonetic information, causing high Gross Pitch Error (GPE) when paired with a vocoder for audio generation, making them unsuitable for end-to-end SpeechLLMs (Sicherman and Adi, 2023; Polyak et al., 2021; Mousavi et al., 2024; Guo et al., 2025).
More recently, a third and more direct paradigm has gained traction: fully supervised training. Given that the primary goal is to capture semantic and phonetic information, this approach directly uses an Automatic Speech Recognition (ASR) objective for supervision. The process involves quantizing the intermediate representations of a powerful ASR encoder and optimizing the model with an ASR loss, ensuring the resulting discrete tokens directly represent linguistic units (Zeng et al., 2025; Du et al., 2024a; b). Subsequently, a downstream vocoder is trained to convert these discrete tokens into mel-spectrograms for speech synthesis. This tokenizer design is foundational to the current state-of-the-art end-to-end SpeechLLMs, underscoring its effectiveness and growing adoption. Interestingly, research has revealed that while the ASR objective targets linguistic content, the resulting tokens retain sufficient extra-phonetic information, such as prosody. This is likely because the ASR encoder implicitly learns to model prosodic features as they serve as valuable auxiliary cues for achieving high transcription accuracy. This retained information allows an integrated LLM to generate highly expressive synthesis and convey complex emotions. Consequently, this design’s ability to support expressive generation has made it a foundational choice for state-of-the-art SpeechLLMs (Zeng et al., 2024; Ding et al., 2025).
@@
R-SPIN (Chang and Glass, 2024) addresses this by learning speaker- and noise-invariant discrete units through a data-efficient self-supervised framework. It extends the speaker-invariant clustering of Spin by using an additional noise-perturbed view of the input and an auxiliary loss that predicts "acoustic pieces," which are phoneme-aligned pseudo-labels, to prevent model collapse and ensure the resulting discrete units represent pure linguistic content . In contrast, NAST (Messica and Adi, 2024) proposes an architecture designed explicitly for robust tokenization, consisting of a predictor, a residual encoder, and a decoder. Its training is governed by a combination of a reconstruction loss, a diversity loss to encourage codebook usage, and a crucial robustness loss that penalizes changes in the predicted token distribution between clean and noise-augmented versions of the same utterance, thereby directly optimizing for token-level stability. Liu et al. (2025) introduce slice-consistency and perturbation-consistency constraints to mitigate discrete representation inconsistency, but their approach targets acoustic tokenizers (rather than semantic tokenizers), which prioritize audio detail reconstruction. Therefore, noise invariance is less meaningful in their context, making their work fundamentally different from ours.
@@ 6 Conclusion
Acknowledgments
This work was supported by National Natural Science Foundation of China (62576010). The corresponding authors are Linhao Zhang and Houfeng Wang.
@@ Acknowledgments
To facilitate reproducibility of our work, we have provided detailed descriptions of the datasets, hyperparameters, and other experimental details used in our study in Section 3 and Appendix B, E, F. Our code and model checkpoint will be released publicly upon acceptance to further support reproducibility and foster future research.
@@ B.1 Training Datasets for StableToken
VCTK (Yamagishi et al., 2019) 44 ASR English
@@
VoiceAssistant-400K (Xie and Wu, 2024) 679 ASR English
AISHELL-1 (Bu et al., 2017) 150 ASR Chinese
@@ B.5 Audio Reconstruction Details
Following the framework of CosyVoice (Du et al., 2024a) and GLM-4-Voice (Zeng et al., 2025), we train a flow matching model to reconstruct audio from speech tokens. The model takes as input the speech token representations and produces Mel spectrograms. Finally, a HiFi-GAN vocoder (Kong et al., 2020) converts the generated Mel spectrograms into the speech waveforms. The datasets for training the flow matching model are listed in Table 10 (excluding our in-house datasets, which comprise both English and Chinese speech data), and the training hyperparameters in Table 11.
B.6 Analysis on Computational Efficiency
B.6.1 Computational Overhead of Different Voter Counts
@@
Table 12: Model parameters and inference FLOPs for different voter counts NN on a 30-second input audio.
@@
B.6.2 Empirical Inference Efficiency Benchmark
From an architectural standpoint, the overhead introduced by the multi-branch design is negligible. The core design of StableToken confines the multi-branch architecture to the lightweight quantization stage only, while the computationally expensive Transformer Encoder remains a single-path process. Since the parallel quantization branches can be executed concurrently with high efficiency on modern hardware, this design introduces almost no additional end-to-end latency compared to conventional single-path tokenizers.
To rigorously evaluate the practical inference efficiency of StableToken, we conducted a comprehensive benchmark test measuring latency, Real-Time Factor (RTF), throughput, and memory footprint. We compared StableToken against a strong baseline, GLM-4-Voice-Tokenizer (Zeng et al., 2025), under identical hardware conditions using both an NVIDIA H20 GPU and an AMD EPYC 9K84 96-Core Processor.
The tests covered various batch sizes and audio durations. The detailed results are summarized in Table 13 (GPU) and Table 14 (CPU).
Table 13: Inference Efficiency and Resource Usage on a single NVIDIA H20 GPU.
Batch Duration (s) Metric StableToken GLM-4-Voice Difference (%)
Size (25 tokens/s) (12.5 tokens/s)
1 1 Latency (ms) 9.11 9.00 +1.22%
RTF 0.0091 0.0090 +1.22%
Throughput (s/s) 109.77 111.11 -1.21%
Memory (MB) 1330 1538 -13.5%
1 30 Latency (ms) 63.17 63.19 -0.03%
RTF 0.0021 0.0021 -0.03%
Throughput (s/s) 474.92 474.79 +0.03%
Memory (MB) 1693 1853 -8.6%
16 10 Latency (ms) 235.96 239.57 -1.5%
RTF 0.0015 0.0015 -1.5%
Throughput (s/s) 678.08 667.87 +1.5%
Memory (MB) 2173 2328 -6.7%
32 60 Latency (ms) 1622.86 1656.49 -2.0%
RTF 0.0008 0.0009 -2.0%
Throughput (s/s) 1183.10 1159.08 +2.1%
Memory (MB) 13978 14230 -1.8%
Table 14: Inference Efficiency and Resource Usage on an AMD EPYC 9K84 96-Core Processor.
Batch Duration (s) Metric StableToken GLM-4-Voice Difference (%)
Size (25 tokens/s) (12.5 tokens/s)
1 1 Latency (ms) 104.27 116.62 -10.6%
RTF 0.1043 0.1166 -10.6%
Throughput (s/s) 9.59 8.57 +11.9%
Memory (MB) 2349 2727 -13.9%
1 30 Latency (ms) 1790.86 1809.86 -1.0%
RTF 0.0597 0.0603 -1.0%
Throughput (s/s) 16.75 16.58 +1.0%
Memory (MB) 2357 2739 -13.9%
16 10 Latency (ms) 5287.85 5687.53 -7.0%
RTF 0.0330 0.0355 -7.0%
Throughput (s/s) 30.26 28.13 +7.6%
Memory (MB) 2566 2982 -14.0%
32 60 Latency (ms) 52021.89 55549.44 -6.4%
RTF 0.0271 0.0289 -6.4%
Throughput (s/s) 36.91 34.56 +6.8%
Memory (MB) 2957 3117 -5.1%
Based on these results, we observe the following: (1) Latency, RTF, and Throughput: The empirical data aligns with our theoretical analysis. The end-to-end latency and RTF of StableToken are nearly identical to the baseline, showing a slight advantage at larger batch sizes (e.g., 2.0% faster at a batch size of 32 on GPU). Throughput matches or slightly exceeds the baseline across all tested scenarios. (2) Memory Footprint: StableToken consistently consumes less memory than the baseline, with savings of up to 13.5% on GPU and 14.0% on CPU. We attribute this efficiency to the use of Lookup-Free Quantization (LFQ), which is inherently more resource-efficient than the Vector Quantization (VQ) employed by the baseline.
In summary, the benchmarks demonstrate that StableToken achieves improved token stability without compromising inference efficiency. Its latency and throughput are competitive with strong baselines, while offering a lower memory footprint.
B.7 Analysis of Token Distribution and Cross-Lingual Efficiency
In this section, we investigate the fundamental properties of the token distribution generated by our multi-branch architecture. Specifically, we analyze the entropy, vocabulary efficiency, and language-specificity of the learned representations. We conduct a statistical analysis on the Chinese and English token distributions to verify whether the architecture effectively learns distinct phonetic representations without introducing significant bias or inefficiency.
Our analysis focuses on two main aspects: (1) Evaluating the efficiency of vocabulary space utilization for individual languages. (2) Assessing the degree of distributional overlap between Chinese and English to confirm the capture of language-specific features.
B.7.1 Vocabulary Utilization and Distribution Entropy
We first analyze the vocabulary usage of StableToken on the LibriSpeech-test (English) and AISHELL (Chinese) datasets. As shown in Table 15, both languages utilize a substantial portion of the available vocabulary. The high entropy values observed for both languages indicate that the token distributions are rich and dispersed, rather than collapsing into a small subset of tokens. This suggests that the tokenizer maintains high representational capacity for both languages.
Table 15: Vocabulary Utilization Statistics on LibriSpeech and AISHELL
Metric English Chinese
Total Tokens 969,440 906,250
Unique Tokens Used 8,139 (99.35%) 7,565 (92.35%)
Token Entropy 12.43 11.68
B.7.2 Cross-Lingual Token Specificity
To assess the language specificity of the learned codebook, we examine the overlap of high-frequency tokens between languages. We identify the top-frequency tokens for one language and calculate their occurrence frequency in the other.
Tables 16 and 17 present these comparisons. The results indicate that tokens with high frequency in one language are extremely rare or non-existent in the other. This disjointed usage pattern demonstrates that the model has successfully learned specialized, language-specific phonetic units, allocating distinct sub-spaces of the codebook to different languages.
Table 16: Frequency Analysis of Top 5 English Tokens
Token ID Freq. (English) Freq. (Chinese) Rank in Chinese
1877 0.4803% 0.0119% 2455
3813 0.3693% 0.00% N/A
3809 0.3377% 0.00% N/A
3812 0.3090% 0.00% N/A
3808 0.2957% 0.0001% 7320
Table 17: Frequency Analysis of Top 5 Chinese Tokens
Token ID Freq. (Chinese) Freq. (English) Rank in English
3810 0.7531% 0.0042% 6276
3811 0.7530% 0.0002% 8045
7910 0.7392% 0.0027% 6850
3815 0.6933% 0.0001% 8104
3234 0.6815% 0.0001% 8085
B.7.3 Quantitative Distributional Divergence
To formally quantify the difference between the token distributions of the two languages, we calculate the Kullback-Leibler (KL) Divergence. As shown in Table 18, the significant non-zero KL Divergence values confirm that the model maintains distinct probability distributions for each language.
Table 18: KL Divergence between Language Token Distributions
Metric Value
KL Divergence (English $\|$ Chinese) 2.81
KL Divergence (Chinese $\|$ English) 2.05
Collectively, these analyses demonstrate that the multi-branch voting mechanism robustly allocates specific regions of the codebook to different languages, capturing unique phonetic characteristics while maintaining efficient vocabulary utilization.
B.8 Long Audio Processing and Boundary Stability
In this section, we clarify the model’s configuration regarding input duration and analyze the stability of tokenization at segment boundaries. This analysis is crucial for understanding the model’s behavior in streaming scenarios or when processing long audio via chunking.
B.8.1 Model Configuration and Segment Length
StableToken inherits the architectural specifications and processing conventions of its backbone, whisper-large-v3. Consequently, the model is configured as follows: (1) Maximum Duration: The tokenizer processes a maximum input audio duration of 30 seconds per inference pass, consistent with the Whisper context window. (2) Training Constraint: During training, all audio segments were constrained to this 30-second window. Audio files exceeding this duration in the training dataset were truncated to the first 30 seconds. (3) Inference Strategy: For audio exceeding 30 seconds, StableToken employs a standard chunking strategy. The input is segmented into 30-second (or smaller) chunks which are processed independently. The resulting token sequences are then concatenated to form the final output.
B.8.2 Analysis of Boundary Stability
A potential concern with chunk-based processing is whether token stability degrades at the boundaries of audio segments (i.e., the beginning and end of a chunk), which could lead to defects when chunks are concatenated together.
To investigate this, we analyzed the distribution of tokenization inconsistencies across different temporal regions of the input audio under Gaussian noise perturbation. We segmented the test audio clips into three regions: (1) Start: The first 15% of the audio duration; (2) Middle: The central 70% of the audio duration; and (3) End: The final 15% of the audio duration. We calculated the Unit Edit Distance (UED) independently for each region. The results are presented in Table 19.
Table 19: Distribution of UED across different temporal regions of audio clips under Gaussian noise.
Region Inconsistencies Ref. Tokens UED % of Total Inconsistencies
Start (0-15%) 16,585 119,770 13.85% 16.29%
Middle (15-85%) 69,333 550,069 12.60% 68.11%
End (85-100%) 15,884 116,016 13.69% 15.60%
Overall 101,802 785,855 12.95% 100.00%
The empirical results indicate that UED is consistent across all regions. The UED in the middle region (12.60%) is only marginally lower than that at the start (13.85%) and end (13.69%). Furthermore, the proportion of total inconsistencies occurring in the start (16.29%) and end (15.60%) regions closely aligns with their respective 15% share of the audio duration.
This uniformity demonstrates that there is no significant degradation in token stability at the segment boundaries. The consensus mechanism employed by StableToken maintains robustness consistently throughout the audio clip, supporting the effectiveness of the chunking strategy for long audio processing.
Appendix C Ablation Study on Architectural Hyperparameters
In this section, we investigate the impact of two key architectural hyperparameters on model performance: the depth at which the quantizer is placed within the encoder, and the ratio of clean to perturbed branches during training. We compare our proposed configuration (L16 + 3:2, which means the quantizer is placed at Layer 16 with a 3:2 clean-to-noisy branch ratio) against three variants: (1) Shallower Quantization (L8 + 3:2): The quantizer is placed at Layer 8; (2) Deeper Quantization (L24 + 3:2): The quantizer is placed at Layer 24; and (3) Reduced Noise Ratio (L16 + 4:1): The ratio of clean to perturbed branches is set to 4:1.
Table 20 summarizes the performance of each model checkpoint at 100k training steps across tokenizer robustness (UED), Automatic Speech Recognition (ASR), and Speech Emotion Recognition (SER) tasks.
Table 20: Ablation study on quantizer placement and perturbation ratio. UED is calculated under Gaussian Noise. LibriSpeech WER is reported for test-clean / test-other. SER is evaluated on the ESD dataset.
Configuration Robustness ASR (WER%, $\downarrow$) SER
(Layer + Clean:Noisy) (UED%, $\downarrow$) LibriSpeech WenetSpeech KeSpeech (Acc%, $\uparrow$)
L16 + 3:2 (Ours) 18.68 2.22 / 5.38 10.91 11.00 67.38
L8 + 3:2 22.05 2.39 / 5.82 11.84 11.24 66.90
L24 + 3:2 13.65 2.52 / 5.96 10.16 11.03 59.51
L16 + 4:1 20.94 2.40 / 5.91 11.00 11.57 62.14
C.1 Effect of Quantizer Placement
The results demonstrate a clear trade-off associated with the quantizer’s depth: (1) Deep Quantization (L24): Placing the quantizer at the deep layer forces it to operate on highly abstract, semantic features. This yields the highest tokenizer robustness (lowest UED) and strong ASR performance on challenging datasets like WenetSpeech. However, at this depth, much of the fine-grained acoustic information (e.g., prosody and timbre) has been abstracted away, resulting in a significant degradation in SER performance (59.51% vs. 67.38%). (2) Shallow Quantization (L8): Operating on shallower features at Layer 8 retains acoustic details but lacks sufficient semantic abstraction. This leads to both poorer tokenizer robustness and degraded ASR performance compared to the deeper configurations. (3) Balanced Configuration (L16): Our proposed placement at Layer 16 strikes an effective balance. It captures sufficiently robust semantic content for high-performance ASR while retaining enough acoustic detail to excel in paralinguistic tasks like SER.
C.2 Effect of Perturbation Ratio
Comparing the 3:2 and 4:1 clean-to-noisy ratios reveals the importance of the training signal difficulty: The 3:2 ratio results in significantly better tokenizer robustness compared to the 4:1 ratio (18.68% vs. 20.94% UED). A higher proportion of perturbed branches creates a more challenging training objective for the consensus mechanism.
This enhanced robustness translates directly to downstream tasks. We hypothesize that the more diverse training signal forces the model to learn representations that are truly invariant to perturbations rather than relying on the clean majority. This richer representation benefits both semantic (ASR) and paralinguistic (SER) performance.
In conclusion, these ablation studies validate the design choices of the proposed model. The combination of Layer 16 placement and a 3:2 noise ratio optimizes the trade-off between semantic robustness and acoustic feature preservation.
Appendix D Baseline Models